OpenStudy (anonymous):

4 log 3 x + log 3 y - 2 log 3 z = log 3 Part 1 [2 points] Is the following equation true? Part 2 [6 points] Use complete sentences to explain the properties used in making your decision

OpenStudy (anonymous):

after the log 3 is (x)sqrt(y)/x^2

jhonyy9 (jhonyy9):

log3(x^4y/z^2)=log3 - so this is right or the base not is 3 ,than what is the base ?

OpenStudy (anonymous):

no after the = sign log 3

jhonyy9 (jhonyy9):

what is the base of log ?

OpenStudy (anonymous):

(x^4)sqrt(y)/(z^2)

jhonyy9 (jhonyy9):

ok i understand now sure - so after the =log3 you will have what i have wrote in my first answer in parantheses

jhonyy9 (jhonyy9):

=log 3 (x^4y)/z^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so wat would be the answers to the 2 questions..

jhonyy9 (jhonyy9):

- so than in 2. points us the propertys of logarithms - log a + log b = log (a*b) - log a - log b = log (a/b) - xlog a = log a^x - hope so much that is understandably sure

jhonyy9 (jhonyy9):

ok ?